# Continuous-Aperture Array (CAPA)-Based Wireless Communications: Capacity Characterization

### v1

@@ Abstract
The capacity limits of continuous-aperture array (CAPA)-based wireless communications are characterized. To this end, an analytically tractable transmission framework is established for both uplink and downlink CAPA systems. Based on this framework, closed-form expressions for the single-user channel capacity are derived. The results are further extended to a multiuser case by characterizing the capacity limits of a two-user channel and proposing the associated capacity-achieving decoding and encoding schemes. 1) For the uplink case, the sum-rate capacity and capacity region, as well as the capacity-achieving detectors, are derived. 2) For the downlink case, the uplink-downlink duality is established by deriving the uplink-to-downlink and downlink-to-uplink transformations under the same power constraint, based on which the optimal power allocation policy and the achieved sum-rate capacity and capacity region are characterized. To gain further insights, several case studies are presented by specializing the derived results into various array structures, including the planar CAPA, linear CAPA, and planar spatially discrete array (SPDA). Numerical results are provided to reveal that: i) the channel capacity achieved by CAPAs converges towards a finite upper bound as the aperture size increases; and ii) CAPAs offer significant capacity gains over the conventional SPDAs.
@@
Channel capacity, continuous aperture array (CAPA), performance analysis.
@@ I Introduction
Multiple antenna technology serves as a cornerstone in the advancement of modern wireless communication systems. At its core lies the principle of leveraging an increased number of antenna elements to enhance spatial degrees of freedom (DoFs) and augment channel capacity [2]. Traditionally, multiple antenna prototypes follow a spatially discrete topological structure, where each antenna is depicted as a distinct point [3]. Motivated by the advantages of incorporating a greater number of antennas, the concept of densely packed antennas has garnered considerable interest in the communication field.
Within a fixed array aperture, reducing the distance between elements enables the addition of more antennas, thus increasing spatial DoFs. This development has led to the emergence of holographic arrays [4, 5]. In such arrays, antenna elements are densely packed in a configuration that is spaced less than half a wavelength apart, resulting in enhanced spectral efficiency [6].
Building on the advancements discussed, this paper introduces a novel concept in antenna design—continuous aperture array (CAPA). CAPA represents a paradigm shift from the conventional spatially discrete array (SPDA), evolving towards an array that forms a spatially continuous electromagnetic (EM) aperture. Essentially, a CAPA can be envisioned as an SPDA comprising an infinite number of antennas with infinitesimal spacing [3, 7]. The concept of a CAPA marks a significant advancement towards maximizing the spatial DoFs available in antenna array configurations, thereby providing substantial gains in system performance. Transitioning from discrete to continuous aperture arrays enhances signal resolution and provides more flexible control over EM wave propagation. Through this innovative approach, CAPA paves the way for next-generation antenna technologies, promising to revolutionize wireless communication infrastructures with its advanced spatial processing capabilities [8].
However, unlike traditional arrays that are modeled using spatially discrete approaches, the modeling of CAPA-based wireless transmission introduces a fundamentally different method grounded in EM field theory. Specifically, while the channels of conventional SPDAs are typically represented through discrete matrices, the spatial response of a CAPA necessitates representation through a continuous operator in Hilbert space [9, 10]. This shift from discrete to continuous modeling is not merely a technical detail but a transformative change in how wireless transmission systems are understood and designed [11]. The use of continuous operators to model the spatial responses of CAPAs signifies a departure from the traditional matrix-based methods suited to SPDAs. This change is crucial because it implies that the entire framework for analyzing and designing transmissions in traditional SPDAs becomes fundamentally inapplicable to CAPAs. Consequently, new conceptual and mathematical frameworks tailored specifically to handle the continuous nature of EM field interactions within CAPAs are necessary [9, 10]. This development is essential for fully exploiting the potential of CAPAs in enhancing system performance.
Currently, research on CAPA remains in a nascent stage and is predominantly focused on single-user scenarios [11, 12, 13, 14, 15]. While these foundational studies have laid the groundwork, exploration into the performance of multiuser CAPA systems is considerably less developed. Their performance, such as channel capacity, has been seldom analyzed, with only [16] presenting some initial results. Specifically, the authors of [16] proposed a sophisticated channel model that utilizes a Green’s function method and leverages a continuous electric current distribution. Their research also extends to the performance analysis of CAPA-based reconfigurable intelligent surfaces (RIS) and simultaneously transmitting and reflecting RISs, which provides new insights into the capabilities of CAPAs in more dynamic settings.
Despite these advancements, a significant gap remains in our understanding of the full potential of CAPAs, particularly regarding their fundamental performance limits. More importantly, an analytically tractable framework for CAPA-based communications is lacking. This framework is essential for advancing both the theoretical understanding and practical deployment of these systems. It would facilitate a deeper comprehension of CAPAs’ capabilities and enable the design of more effective and efficient communication strategies tailored to their unique properties, especially in multiuser scenarios where complexity and demand for system efficiency are significantly higher.
Motivated by these knowledge gaps, this paper proposes a novel analytical framework to address the continuous nature of EM fields within CAPAs, providing a robust foundation for modeling and understanding CAPA-based communications. Using this framework, the channel capacity for both downlink and uplink CAPA-based communications is characterized. The main contributions are summarized as follows.
@@
  • We propose a transmission framework for uplink and downlink CAPA communications using EM field theories, where continuous operators model both signals and spatial responses. Based on this framework, we design optimal detectors and source current distributions to achieve the single-user channel capacity for uplink and downlink CAPA communications, and derive closed-form expressions for the corresponding channel capacity.
@@
  • We then extend our analysis to multiuser CAPA systems by studying a two-user case. For uplink CAPA communications, we propose using successive decoding to achieve capacity and derive closed-form expressions for the optimal detectors of each user. Based on this, we derive closed-form expressions for the per-user rate and the sum-rate capacity, and characterize the capacity region. For downlink CAPA communications, we establish the uplink-downlink duality by presenting the uplink-to-downlink and downlink-to-uplink transformations under the same sum power constraint and derive closed-form expressions for the capacity-achieving source current distributions. Additionally, we characterize the downlink capacity limits by deriving the per-user rate, sum-rate capacity, and capacity region.
@@
  • To unveil further system insights, we present several case studies exploring the line-of-sight (LoS) channel model. The derived results are specialized to various array structures, including the planar CAPA, linear CAPA, and planar SPDA. We derive closed-form expressions for the sum-rate capacity in each case and perform asymptotic analyses by setting the aperture size to infinity.
@@
  • We present numerical results illustrating that the capacity achieved by the CAPA increases with aperture size and converges to a finite upper limit. Our results demonstrate that, compared to CAPA, the capacity of the conventional SPDA is limited by its array occupation ratio. Moreover, CAPA offers significant capacity enhancements over conventional SPDA, whose capacity converges to that of CAPA as the array occupation ratio approaches one.
@@
The remainder of this paper is organized as follows. Section II introduces the transmission framework for CAPA-based communications. Sections III and IV delve into the capacity analysis and the capacity-achieving schemes for single-user and multiuser CAPA systems, respectively. Section V provides case studies for several specific array structures. Section VI presents numerical results to demonstrate the superior performance of CAPAs. Finally, Section VII concludes the paper.
@@ II-A1 Uplink Communications
${\mathsf{E}}_{{\mathsf{ul}},k}(\mathbf{r})$ $=\int_{{\mathcal{A}}_{k}}{\mathsf{H}}(\mathbf{r},{\mathbf{s}}){x}_{{\mathsf{ul}},k}({\mathbf{s}}){\rm{d}}{\mathbf{s}}$ (1a)
$=s_{{\mathsf{ul}},k}\int_{{\mathcal{A}}_{k}}{\mathsf{H}}(\mathbf{r},{\mathbf{s}}){\mathsf{J}}_{{\mathsf{ul}},k}({\mathbf{s}}){\rm{d}}{\mathbf{s}},$ (1b)
@@ II-A2 Downlink Communications
${x}_{{\mathsf{dl}}}({\mathbf{r}})=\sum\nolimits_{k=1}^{K}{\mathsf{J}}_{{\mathsf{dl}},k}({\mathbf{r}})s_{{\mathsf{dl}},k}$ (4)
@@
Since the aperture size of each user is generally negligible with respect to both the propagation distance and the BS aperture size, the variations of the electric field and channel response along the receive aperture can be omitted. This yields ${\mathsf{E}}_{{\mathsf{dl}},k}(\mathbf{s})\approx{\mathsf{E}}_{{\mathsf{dl}},k}(\mathbf{s}_{k})\triangleq{\mathsf{E}}_{{\mathsf{dl}},k}$ and ${\mathsf{H}}(\mathbf{s},{\mathbf{r}})\approx{\mathsf{H}}(\mathbf{s}_{k},{\mathbf{r}})={\mathsf{H}}({\mathbf{r}},\mathbf{s}_{k})=\mathsf{H}_{k}(\mathbf{r})$ for $\mathbf{s}\in{\mathcal{A}}_{k}$. Therefore, we obtain
@@ II-B Channel Model
To facilitate theoretical investigations into fundamental performance limits and asymptotic behaviors, we focus our discussion on LoS channels. Under this consideration, ${\mathsf{H}}(\mathbf{r},{\mathbf{s}})$ is modeled as follows [19]:
@@
${\mathsf{H}}^{\mathsf{em}}(\mathbf{r},{\mathbf{s}})\triangleq\frac{{\rm{j}}k_{0}\eta{\rm{e}}^{-{\rm{j}}k_{0}\lVert{\mathbf{r}}-{\mathbf{s}}\rVert}}{4\pi\lVert{\mathbf{r}}-{\mathbf{s}}\rVert}\left(1+\frac{{\rm{j}}k_{0}^{-1}}{\lVert{\mathbf{r}}-{\mathbf{s}}\rVert}-\frac{k_{0}^{-2}}{\lVert{\mathbf{r}}-{\mathbf{s}}\rVert^{2}}\right)$ (9)
models the influence of free-space EM propagation, where $\eta=120\pi$ is the impedance of free space, and $k_{0}=\frac{2\pi}{\lambda}$ with $\lambda$ being the wavelength denotes the wavenumber. Particularly, ${\mathsf{H}}^{\mathsf{em}}(\mathbf{r},{\mathbf{s}})$ comprises three terms: the first term corresponds to the radiating near-field and far-field regions, while the remaining two terms correspond to the reactive near-field region. The squared magnitude of the parenthesis in (9) is $1-\frac{k_{0}^{-2}}{\lVert{\mathbf{r}}-{\mathbf{s}}\rVert^{2}}+\frac{k_{0}^{-4}}{\lVert{\mathbf{r}}-{\mathbf{s}}\rVert^{4}}$, which equals $0.97$ at distance $\lVert{\mathbf{r}}-{\mathbf{s}}\rVert=\lambda$ [19]. Hence, when considering practical systems with $\lVert{\mathbf{r}}-{\mathbf{s}}\rVert\gg\lambda$, the last two terms in (9) can be neglected. As a result, (9) is approximated as follows:
${\mathsf{H}}^{\mathsf{em}}(\mathbf{r},{\mathbf{s}})\approx\frac{{\rm{j}}k_{0}\eta}{4\pi\lVert{\mathbf{r}}-{\mathbf{s}}\rVert}{\rm{e}}^{-{\rm{j}}k_{0}\lVert{\mathbf{r}}-{\mathbf{s}}\rVert}.$ (10)
III Single-User Case

### v2

@@ Abstract
The capacity limits of continuous-aperture array (CAPA)-based wireless communications are characterized. To this end, an analytically tractable transmission framework is established for both uplink and downlink CAPA systems. Based on this framework, closed-form expressions for the single-user channel capacity are derived. The results are further extended to a multiuser case by characterizing the capacity limits of a two-user channel and proposing the associated capacity-achieving decoding and encoding schemes. In the uplink case, the capacity-achieving detectors and sum-rate capacity are derived, and the capacity region is characterized. In the downlink case, the uplink-downlink duality is established by deriving the uplink-to-downlink and downlink-to-uplink transformations under the same power constraint, based on which the optimal source current distributions and the achieved sum-rate capacity and capacity region are characterized. For comparison, the uplink and downlink sum-rates achieved by the linear zero-forcing scheme are also analyzed. To gain further insights, several case studies are presented by specializing the derived results into various array structures, including the planar CAPA, linear CAPA, and planar spatially discrete array (SPDA). Numerical results are provided to reveal that the channel capacity achieved by CAPAs converges towards a finite upper bound as the aperture size increases; and CAPAs offer superior capacity over the conventional SPDAs.
@@
Channel capacity, continuous-aperture array (CAPA), downlink, performance analysis, uplink.
@@ I Introduction
Multiple-antenna technology is considered one of the fundamental building blocks of modern wireless communication systems. At its core lies the principle of utilizing an increased number of antenna elements and expanded array apertures to enhance spatial degrees of freedom (DoFs) and array gains, ultimately improving channel capacity [2, 3]. In this context, technologies such as multiple-input multiple-output (MIMO) and massive MIMO have played a crucial role in advancing cellular networks from the third generation (3G) to the fourth (4G), fifth (5G), and beyond [4, 5, 6].
In general, the number of spatial DoFs and array gains scale with the number of antenna elements and the array aperture size. Motivated by this principle, various advanced MIMO variants and novel array architectures have been proposed. Examples include holographic MIMO [7], gigantic MIMO [8], extremely large aperture arrays [9], reconfigurable intelligent surfaces [10], dynamic metasurface antennas [11], and fluid [12], movable [13], or pinching [14] antennas. Extensive theoretical analyses and experimental studies have demonstrated the substantial performance gains achievable through these emerging multiple-antenna technologies. Despite their diverse architectures, these novel multiple-antenna systems share a common evolutionary trend: larger aperture sizes, denser antenna configurations, higher operating frequencies, and more flexible structures. This progression paves the way for the development of an (approximately) continuous electromagnetic (EM) aperture [15, 16, 17], commonly referred to as a continuous-aperture array (CAPA) [18, 19, 20], which is to be focused on in this paper.
A CAPA functions as a single, large-aperture antenna with a continuous current distribution, which comprises a (virtually) infinite number of radiating elements integrated with electronic circuits and driven by a limited number of radio-frequency (RF) chains. Unlike conventional spatially discrete arrays (SPDAs), a CAPA fully exploits the entire aperture surface and allows for precise control over the current distribution [20]. This capability significantly enhances spatial DoFs and array gains, as demonstrated in various research endeavors; see Section I-A2 for more relevant references. However, the continuous nature of EM field interactions in CAPAs necessitates a fundamental shift in system modeling. Instead of the matrix-based models used for SPDAs, CAPA-based systems require an EM-theoretic approach, leading to a continuous Hilbert–Schmidt operator-based integral framework [21, 22, 23, 24, 25, 26, 27]. This transition from discrete to continuous modeling is not just a mathematical refinement but a paradigm shift in the design and analysis of wireless transmission systems. To fully unlock the potential of CAPAs, a novel conceptual framework tailored to their unique characteristics is essential.
I-A Prior Works
I-A1 Hardware Implementation of CAPAs
The development of semi-CAPA and CAPA architectures has a long history in antenna design research, dating back to pioneering contributions in the 1960s [28, 29]. With recent advancements in materials science and array fabrication, researchers have successfully developed CAPA prototypes based on designs such as metasurface-based leaky-wave antennas [30, 31], optically driven tightly coupled arrays [32], and interdigital transducer-based grating antennas [33]; see [34, 20] for further details. Some of these prototypes have even reached commercialization, with early test reports demonstrating their significant potential for enhancing coverage and throughput in practical wireless networks [35, 36].
It is worth noting that most CAPA prototypes available today are semi-continuous; however, they are effectively treated as continuous antennas, as emphasized by their inventors [35, 36]. Achieving a true CAPA would require modulating phase and amplitude at each aperture point in an analog manner. One possible realization involves using a metasurface composed of materials such as photographic emulsion, which, when illuminated by a tailored wave, could enable unique phase and amplitude modulation at each point. This approach would ultimately generate the desired continuous current distribution when a reference wave propagates over the surface [4, 34]. Such advancements align with the objectives of leading enterprises, which are striving for an advanced level of spatial continuity in antenna design.
Besides the aforementioned array architectures, several key challenges must be carefully addressed to ensure the practicality and feasibility of CAPA and semi-CAPA systems. Some of the most critical challenges include:
  • EM Coupling: When antennas are densely or continuously packed, mutual coupling (MC) effects become significant. MC occurs where EM waves transmitted by one antenna are absorbed by adjacent antennas, which alters their circuitry and distorting the overall spatial channel [37]. In a CAPA or an ultra-dense semi-CAPA, EM coupling can lead to severe performance degradation, increased interference, and complex interactions that are difficult to predict and mitigate.
  • Complex Feed Network: A CAPA or an ultra-dense semi-CAPA requires a highly complex feed network to properly excite the aperture or individual elements. Designing such a network for a continuous surface or a large number of densely packed elements introduces substantial complexity, increases insertion losses, and raises overall power consumption [38, 39].
  • Material Limitations: The materials used to construct and sustain a CAPA or an ultra-dense semi-CAPA must meet stringent physical requirements. Challenges such as thermal management, mechanical durability, and environmental susceptibility (e.g., humidity, temperature variations, and material aging) impose additional constraints on long-term reliability and scalability [38, 39].
Taken together, these challenges indicate that realizing a fully functional CAPA or a densely packed semi-CAPA remains elusive, even with recent advancements in materials and fabrication techniques. Addressing these issues requires collaborative efforts from both the communication and antenna communities.
I-A2 Theoretical Design and Analysis of CAPAs
Although research on the hardware implementation of an ideal CAPA is still in its infancy, increasing efforts have been devoted to the theoretical design and analysis of CAPA- or semi-CAPA-based communications. Many of these studies rely on simplified models, such as neglecting the effects of EM coupling and considering narrowband scenarios. One of the most widely studied topics is the number of EM DoFs [6], which has been investigated using various analytical tools, including diffraction theory [40], Nyquist-Shannon sampling theorem [41, 21], and Landau’s theorem [42]. Beyond DoFs, the array gain achieved by CAPAs has also been analyzed in [43]. These initial studies collectively highlight the advantages of CAPAs over traditional SPDAs.
In addition, significant research attention has been given to characterizing the channel capacity of CAPA-based point-to-point communication systems. These studies leverage Green’s function-based models to describe the EM channel and apply Hilbert operator theory to derive capacity expressions under various propagation environments [23, 24, 22, 26, 25, 27]. Recently, this research has been extended to scenarios where a CAPA-based reconfigurable surface assists communications between SPDAs [44]. Another key research direction focuses on continuous beamforming design for CAPA-based communications. One prevalent approach relies on discretization-based methods, such as Fourier series expansion, to approximate continuous signals and simplify the associated optimization problems [17, 19]. Additionally, calculus of variations has been applied to design low-complexity beamforming strategies [45] and explore optimal beamforming structures [46, 47].
As mentioned earlier, these research developments have relied on simplified system models that primarily consider the spatially continuous nature of CAPAs while omitting other crucial factors, such as the impact of MC and the challenges posed by wideband scenarios. To address these limitations, the authors in [48] proposed an EM-based array manifold that accounts for complex physical interactions, including MC. This framework enables the modeling of arbitrary antenna configurations, including CAPAs, with greater physical accuracy.
Regarding wideband transmission, the work in [49] introduced a small-scale fading model for CAPA-based communications using $k$-space (wavenumber domain) analysis. This approach provides a wavenumber (spatial-frequency) representation of how the aperture is excited, which illustrates how energy is distributed across different propagation directions (angles) at each frequency. Such insights can be instrumental in tackling spatial-wideband effects in CAPA-based wideband communications. Another research direction involves circuit-based modeling. A recent example in [50] employs multi-port circuit theory combined with Chu’s antenna model to investigate the impact of MC on the wideband behavior of ultra-dense arrays.
I-B Motivation and Contributions
Despite these advantages, the fundamental theoretical limits of CAPA-based communications remain largely unexplored. Building on existing research in CAPA-based single-user communications, this paper takes a step forward by investigating the capacity limits of CAPA-based multiuser systems and establishing an analytically tractable framework for multiuser CAPA communications. Specifically, we consider a scenario where a CAPA serves multiple users equipped with traditional single-antenna devices. For simplicity, we focus on a narrowband model, while extending the analysis to wideband scenarios remains an important avenue for future research. The main contributions are summarized as follows.
@@ I-B Motivation and Contributions
  • We propose a transmission framework for uplink and downlink CAPA communications using EM field theories, where continuous operators model both signals and spatial responses. Based on this framework, we design optimal detectors and source current distributions to achieve the single-user channel capacity and derive closed-form expressions for the corresponding capacity.
@@
  • We extend our analysis to multiuser CAPA systems by studying a two-user case. For uplink communications, we propose using successive decoding to achieve capacity and derive closed-form expressions for the optimal detectors of each user. Based on this, we derive closed-form expressions for the per-user rate and the sum-rate capacity, and characterize the capacity region. For downlink communications, we establish the uplink-downlink duality by presenting the uplink-to-downlink and downlink-to-uplink transformations under the same sum power constraint and derive closed-form expressions for the capacity-achieving source current distributions. Additionally, we characterize the downlink capacity limits by deriving the per-user rate, sum-rate capacity, and capacity region. For comparison, we also analyze the uplink and downlink sum-rates achieved using the linear zero-forcing (ZF) detector and precoder, respectively.
@@
  • To unveil further insights, we present several case studies exploring the line-of-sight (LoS) channel model. The derived results are specialized to various array structures, including the planar CAPA, linear CAPA, and planar SPDA. We derive closed-form expressions for the sum-rate capacity in each case and perform asymptotic analyses by setting the aperture size to infinity.
@@
  • We present numerical results illustrating that the capacity achieved by the CAPA increases with aperture size and converges to a finite upper limit. Our results demonstrate that, compared to CAPA, the capacity of the conventional SPDA is limited by its array occupation ratio. Moreover, CAPA can theoretically achieve greater capacity than conventional SPDA, whose capacity converges to that of CAPA as the array occupation ratio approaches one.
@@
The remainder of this paper is organized as follows. Section II introduces the transmission framework for CAPA-based communications. Sections III and IV delve into the capacity analysis and the capacity-achieving schemes for uplink and downlink transmission, respectively. Section V provides case studies for several specific array structures. Section VI presents numerical results to demonstrate the superior performance of CAPAs. Finally, Section VII concludes the paper.
@@ II-A1 Uplink Communications
${\mathsf{E}}_{{\mathsf{ul}},k}(\mathbf{r})$ $=\int_{{\mathcal{A}}_{k}}{\mathsf{H}}(\mathbf{r},{\mathbf{s}}){x}_{{\mathsf{ul}},k}({\mathbf{s}}){\rm{d}}{\mathbf{s}},$ (1)
@@ II-A2 Downlink Communications
${x}_{{\mathsf{dl}}}({\mathbf{r}})=\sum\nolimits_{k=1}^{K}{\mathsf{J}}_{{\mathsf{dl}},k}({\mathbf{r}})s_{{\mathsf{dl}},k},$ (4)
@@
Since the variations of the electric field and channel response along the receive aperture are negligible, we have ${\mathsf{E}}_{{\mathsf{dl}},k}(\mathbf{s})\approx{\mathsf{E}}_{{\mathsf{dl}},k}(\mathbf{s}_{k})\triangleq{\mathsf{E}}_{{\mathsf{dl}},k}$ and ${\mathsf{H}}(\mathbf{s},{\mathbf{r}})\approx{\mathsf{H}}(\mathbf{s}_{k},{\mathbf{r}})$ for $\mathbf{s}\in{\mathcal{A}}_{k}$. By further assuming channel reciprocity between the uplink and downlink, we have ${\mathsf{H}}(\mathbf{s}_{k},{\mathbf{r}})={\mathsf{H}}({\mathbf{r}},\mathbf{s}_{k})=\mathsf{H}_{k}(\mathbf{r})$, which gives
@@ II-B Channel Model
To facilitate theoretical investigations into fundamental performance limits, we focus our discussion on LoS channels. In this case, ${\mathsf{H}}(\mathbf{r},{\mathbf{s}})$ is modeled as follows [43]:
@@
${\mathsf{H}}^{\mathsf{em}}(\mathbf{r},{\mathbf{s}})\triangleq\frac{{\rm{j}}k_{0}\eta{\rm{e}}^{-{\rm{j}}k_{0}\lVert{\mathbf{r}}-{\mathbf{s}}\rVert}}{4\pi\lVert{\mathbf{r}}-{\mathbf{s}}\rVert}$ (9)
models the influence of free-space EM propagation [43], where $\eta=120\pi\,\Omega$ is the impedance of free space, and $k_{0}=\frac{2\pi}{\lambda}$ with $\lambda$ being the wavelength denotes the wavenumber.
If we consider a rich-scattering environment, ${\mathsf{H}}(\mathbf{r},{\mathbf{s}})$ can be modeled as a stochastic process [7, 49]. Additionally, the channel model in (8) neglects the impact of MC. If MC is taken into account, we could model ${\mathsf{H}}(\mathbf{r},{\mathbf{s}})$ as follows:
${\mathsf{H}}(\mathbf{r},{\mathbf{s}})=\int_{{\mathcal{A}}}\mathsf{Z}_{\mathsf{mc}}(\mathbf{r},\mathbf{r}^{\prime}){\mathsf{H}}_{\mathsf{nmc}}(\mathbf{r}^{\prime},{\mathbf{s}}){\rm{d}}\mathbf{r}^{\prime},$ (10)
which generalizes the commonly used linear MC model [37]. Here, the operator $\mathsf{Z}_{\mathsf{mc}}(\mathbf{r},\mathbf{r}^{\prime})$ captures the MC effects within the CAPA, while ${\mathsf{H}}_{\mathsf{nmc}}(\mathbf{r}^{\prime},{\mathbf{s}})$ represents the spatial channel response without considering MC. Modeling $\mathsf{Z}_{\mathsf{mc}}(\mathbf{r},\mathbf{r}^{\prime})$ requires circuit-theoretic knowledge [50, 37], which is beyond the scope of this paper. A detailed study of MC effects on CAPA-based communications is left for future work, although some initial simulation results are provided in Section VI.
For the rest of this paper, we adopt the LoS channel model as an illustrative example. However, it is important to emphasize that our proposed approaches are not constrained to any specific channel type, nor do they rely on assumptions about MC. Since all the derived results are based on continuous operator formulations, they remain applicable in a wide range of channel environments.
III Uplink Capacity
This section characterizes the channel capacity for CAPA-based uplink communications. We commence with the single-user case, and then proceed to explore the two-user case, while extensions to scenarios with an arbitrary number of users will be addressed in future work.
III-A Single-User Case